The balance sheet for the Delphine, Xavier, and Olivier partnership follows: Cash Noncash assets 71,520 Liabilities 48,000 132,000 60,960 56,000 38,560 203,520 Delphine, capital Xavier, capital Olivier, capital Total liabilities and capital Total assets 203,520 Delphine, Xavier, and Olivier share profits and losses in the ratio of 3:4:3, respectively. The partners have agreed to terminate the business and estimate that $15,200 in liquidation expenses will be incurred. Which partner should receive the cash distribution?
Answer #1

Assuming all non-cash assets and liabilities can be sold/paid at their balance sheet value.

Total Cash available from Assets = Cash + Non-cash assets

= 71,520 + 132,000

= 203,520

Total available cash for distribution = Cash available from assets - Liabilities - Liquidation expenses

= 203,520 - 48,000 - 15,200

= 140,320

Distribution among partners:

Partner Name Profit/Loss Distribution ratio Distribution % Distribution $ Capital $
Delphine 3 3/3+4+3 = 3/10 = 30%

30%*140320

= 42,096

60,960
Xavier 4 4/3+4+3 = 4/10 = 40%

40%*140320

= 56,128

56,000
Olivier 3 3/3+4+3 = 3/10 = 30%

30%*140320

= 42,096

38,560

Since, Delphine shall receive only $42,096 in the liquidation but her exposure (capital) is more, that is, $60,960, she shall be eligible to receive a cash distribution to the extent 60960 - 42096 = $18,864